# RIFT: Closed-Loop RL Fine-Tuning for Realistic and Controllable Traffic Simulation

### v2

@@ 5.4 Ablation Study (Q3)
In exploring the design choices within RIFT, we conduct a series of ablation studies focusing on four key aspects: the choice of policy optimization clip (Dual-Clip vs. PPO-Clip), the fine-tuning module (Scoring Head vs. All Head), the choice of KL regularization (w/ KL vs. w/o KL), and the reward design (Normal vs. Aggressive). All experiments are conducted under consistent settings, and results are summarized in Table˜3.
Dual-Clip vs. PPO-Clip. We first evaluate the effect of the dual-clip mechanism by replacing it with the standard PPO-Clip (w/ PPO-Clip). This variant leads to more conservative and less realistic behaviors, as evidenced by reduced driving efficiency and degraded kinematic metrics.
Scoring Head vs. All Head. Next, we examine the impact of the fine-tuning module on performance by comparing RIFT fine-tuned using only the scoring head with a variant fine-tuned on all heads (w/ All Head). The latter induces a slight degradation across metrics, primarily due to the disruption of trajectory-level realism learned during pre-training.
w/ KL vs. w/o KL. Then, we ablate the KL term that constrains divergence from the IL pre-trained policy. As shown in Table˜3, removing this constraint improves controllability without degrading realism. This validates our decoupled design: instead of enforcing distributional consistency via KL, we freeze the generation head and fine-tune only the scoring head. This separation enables flexible trajectory selection while preserving the trajectory-level realism inherited from pre-training.
Normal vs. Aggressive. Finally, we explore the effect of different reward preferences on controllability by adopting a more aggressive reward focused on driving efficiency (w/ Aggressive). This variant improves driving efficiency but increases collision frequency and off-road rate, highlighting the intended trade-off and demonstrating that RIFT supports flexible reward design while maintaining style-level controllability. Additional qualitative results regarding controllability are presented in Appendix D.1.
@@ 6 Conclusion
In this work, we propose a dual-stage AV-centric simulation framework that conducts open-loop IL pre-training in a data-driven simulator to capture trajectory-level realism and route-level controllability, followed by closed-loop RL fine-tuning in a physics-based simulator to address covariate shift and enhance style-level controllability. During fine-tuning, we introduce RIFT, a novel RL fine-tuning strategy that evaluates all candidate modalities using group-relative optimization combined with a dual-clip surrogate objective, thereby enhancing style-level controllability and mitigating covariate shift, while preserving the trajectory-level realism and route-level controllability established in IL pre-training. Extensive experiments demonstrate that RIFT generates scenarios with superior realism and controllability, effectively revealing the limitations of modern AV systems and further bridging the gap between traffic simulation and reliable closed-loop evaluation.
Limitations and Future Work. In the current framework, the generation head is frozen during closed-loop fine-tuning, and its reliability solely stems from the robust trajectory generation capability learned during IL pre-training. However, without reliable expert demonstrations, the realism and robustness of generated trajectories cannot be further improved during closed-loop fine-tuning. This limitation highlights a key avenue for future research: developing methods—potentially leveraging RL or other self-improvement paradigms—that can enhance the quality of trajectory generation without relying on expert demonstrations.
@@ Appendix A Theoretical Analysis
This section thoroughly analyzes RIFT under a binary reward setting ($r\in\{0,1\}$) without any KL regularization. The dual-clip surrogate guarantees stable gradients and a globally convergent update in success-probability space.
A.1 Problem Setup
For a given state $s$, let $\pi_{\theta}(\tau\mid s)$ denote the probability of trajectory $\tau$ under policy parameters $\theta$. For simplicity, we use $\tau$ to denote both the rollout and the candidate trajectory, as there exists a fixed mapping between them. Define the success probability
$p_{\theta}(s)\;:=\;\Pr_{\tau\sim\pi_{\theta}(\cdot\mid s)}\bigl{(}r(\tau,s)=1\bigr{)},$
and write $p:=p_{\theta_{\text{old}}}(s)\in(0,1)$ for the current policy. Since $r\in\{0,1\}$, the mean and variance of the Bernoulli random variable are given by:
$\mathbb{E}_{\tau\sim\pi_{\theta_{\text{old}}}}[r(\tau,s)]=p,\qquad\mathsf{Var}_{\tau\sim\pi_{\theta_{\text{old}}}}[r(\tau,s)]=p(1-p).$
The (group-)relative advantage shao2024grpo in Equation˜6 therefore reduces to
$\hat{A}(\tau,s)=\begin{cases}\sqrt{\tfrac{1-p}{p}},&r(\tau,s)=1,\\[6.0pt] -\sqrt{\tfrac{p}{1-p}},&r(\tau,s)=0.\end{cases}$ (9)
Dual-Clip Surrogate. RIFT discards the KL term entirely and maximizes
$f_{\epsilon}^{\mathrm{RIFT}}(\rho_{i},\hat{A}_{i})=\max\!\Bigl{(}\min\!\bigl{(}\rho_{i}\hat{A}_{i},\,\operatorname{clip}(\rho_{i},1-\epsilon,1+\epsilon)\hat{A}_{i}\bigr{)},c\,\hat{A}_{i}\Bigr{)},$ (10)
where $\rho_{i}=\pi_{\theta}(\tau_{i}\mid s)/\pi_{\theta_{\text{old}}}(\tau_{i}\mid s)$ is the trajectory-level importance sampling ratio, $\epsilon>0$ is the clipping range, and $c>1$ caps the most negative value.
Remark A.1.
If $\hat{A}_{i}<0$ and $\rho_{i}\gg 1$, the naïve PPO/GRPO objective becomes a large negative value, producing unstable gradients. Because
$f_{\epsilon}^{\mathrm{RIFT}}(\rho_{i},\hat{A}_{i})\;\geq\;c\,\hat{A}_{i},$ (11)
the surrogate is never more negative than $c\hat{A}_{i}$, so
$\bigl{|}\nabla_{\theta}\mathbb{E}[f_{\epsilon}^{\mathrm{RIFT}}]\bigr{|}\;\leq\;\bigl{|}\nabla_{\theta}\mathbb{E}[\rho_{i}\hat{A}_{i}]\bigr{|}.$ (12)
Hence, dual-clip reduces the worst-case gradient magnitude, ensuring stability.
A.2 Iterative Update in the Success-Probability Space
Let $\theta_{\text{new}}\in\arg\max_{\theta}\mathbb{E}_{\tau\sim\pi_{\theta_{\text{old}}}}\!\bigl{[}f_{\epsilon}^{\mathrm{RIFT}}(\rho_{i},\hat{A}_{i})\bigr{]}$ and denote the corresponding success probability by $p_{\text{new}}:=p_{\theta_{\text{new}}}(s)$. The update may therefore be viewed as a mapping
$p_{n}=h_{\epsilon}(p_{n-1}),$ (13)
with $p_{n-1}=p_{\theta_{\text{old}}}(s)$ and $p_{n}=p_{\text{new}}$. We prove:
  1. (i)
    $h_{\epsilon}:[0,1]\to(0,1)$ is well-defined and continuous;
  2. (ii)
    with suitable $(c,\epsilon)$ or a sufficiently small step size, $h_{\epsilon}$ is a contraction, yielding a unique fixed point $p^{\star}\in(0,1)$ and global convergence $p_{n}\to p^{\star}$.
A.3 Existence of a Fixed Point
Lemma A.2 (Continuity and Interval Mapping).
For every $p\in[0,1]$ the updated probability $h_{\epsilon}(p)$ lies in $(0,1)$, and $h_{\epsilon}$ is continuous on $[0,1]$.
Proposition A.3 (Existence).
At least one fixed point $p^{\star}\in(0,1)$ satisfies $p^{\star}=h_{\epsilon}(p^{\star})$.
A.4 Uniqueness and Global Convergence
Bounded derivatives. For $\hat{A}_{i}\geq 0$, $|\partial f_{\epsilon}^{\mathrm{RIFT}}/\partial\rho_{i}|\leq(1+\epsilon)|\hat{A}_{i}|;$ for $\hat{A}_{i}<0$ it is bounded by $|\hat{A}_{i}|$. Consequently, the dependence of $p_{\text{new}}$ on $p$ admits a global Lipschitz constant $L(c,\epsilon)$.
Contraction. If $L(c,\epsilon)<1$ or a step size $\eta<1/L(c,\epsilon)$ is used, $h_{\epsilon}$ is a strict contraction on $(0,1)$.
Theorem A.4 (Convergence of RIFT).
Assume $c>1$ and $\epsilon>0$. If $|h_{\epsilon}(p)-h_{\epsilon}(p^{\prime})|\leq\lambda|p-p^{\prime}|$ for some $\lambda<1$ and all $p,p^{\prime}\in(0,1)$—for example, by choosing a step size $\eta<1/L(c,\epsilon)$—then
  1. (a)
    $h_{\epsilon}$ possesses a unique fixed point $p^{\star}\in(0,1)$;
  2. (b)
    any sequence $p_{n}=h_{\epsilon}(p_{n-1})$ converges geometrically: $|p_{n}-p^{\star}|\leq\lambda^{n}|p_{0}-p^{\star}|$;
  3. (c)
    The dual-clip term eliminates gradient explosions caused by negative advantages, guaranteeing stable learning.
A.5 Detailed Proofs
Throughout this subsection, we reuse the notation introduced in Section˜A.4:
$A^{+}(p)=\sqrt{\frac{1-p}{p}},\quad A^{-}(p)=-\sqrt{\frac{p}{1-p}},\quad\rho^{+}=1+\epsilon,\;\rho^{-}=1-\epsilon,\quad p\in(0,1),\ \epsilon>0,\ c>1.$
Proof of Lemma˜A.2.
The proof consists of three parts.
Boundary points. If $p=0$, every trajectory under the old policy fails and $A^{+}(0^{+})=\infty$. For any candidate new policy with success probability $q\in(0,1]$,
$\mathcal{J}(q)=q\,\rho^{+}A^{+}(0^{+})+(1-q)\,c\,A^{-}(0^{+})=(1+\epsilon)\,q\cdot\infty>0,\qquad\mathcal{J}(0)=0.$
Hence $h_{\epsilon}(0)=\arg\max_{q}\mathcal{J}(q)>0$. The case $p=1$ is symmetric, yielding $h_{\epsilon}(1)<1$.
Interior points. Let $p\!\in\!(0,1)$ and denote the step in parameter space by $\Delta\theta$. Smoothness of $\theta\!\mapsto\!\pi_{\theta}$ implies $\bigl{|}\log\rho_{i}\bigr{|}\!\leq\!K\|\Delta\theta\|$ for some $K<\infty$. Because $f_{\epsilon}^{\mathrm{RIFT}}$ is $(1+\epsilon)$-Lipschitz in $\rho_{i}$, the expected surrogate, cannot change by more than $K(1+\epsilon)\|\Delta\theta\|$ in one update; choosing $\|\Delta\theta\|<\!(K(1+\epsilon))^{-1}$ keeps $h_{\epsilon}(p)$ strictly inside $(0,1)$.
Continuity. Set $\Psi(\theta,p)=\mathbb{E}_{\tau\sim\pi_{\theta}}\![f_{\epsilon}^{\mathrm{RIFT}}]$. Joint continuity of $\Psi$ and compactness of the trajectory space (after clipping) satisfies the hypotheses of Berge’s maximum theorem, so the maximizer correspondence $\arg\max_{\theta}\Psi(\theta,p)$ is upper-hemicontinuous; the composition yields continuity of $h_{\epsilon}$ on $[0,1]$. ∎
Proof of Proposition˜A.3.
Define $g(p)=h_{\epsilon}(p)-p$. Lemma˜A.2 gives $g(0)>0$ and $g(1)<0$, while Lemma˜A.2 also guarantees the continuity of $g$. The Intermediate Value Theorem therefore yields a root $p^{\star}\in(0,1)$ with $p^{\star}=h_{\epsilon}(p^{\star})$. ∎
Proof of Theorem˜A.4.
The proof proceeds in four steps.
Step 1 – maximized surrogate. Substituting $\rho^{+},\rho^{-}$ into the surrogate gives
$\Phi(p)=p\,\rho^{+}A^{+}(p)+(1-p)\,\rho^{-}A^{-}(p)=2\epsilon\,\sqrt{p(1-p)}.$ (14)
Step 2 – Lipschitz bound.
$\Phi^{\prime}(p)=\epsilon\,\frac{1-2p}{\sqrt{p(1-p)}},\quad|\Phi^{\prime}(p)|\leq\frac{\epsilon}{\sqrt{p(1-p)}}.$ (15)
Fix $\delta\in(0,\tfrac{1}{2})$ such that the first iterate satisfies $h_{\epsilon}(p_{0})\in[\delta,1-\delta]$ (Lemma˜A.2 ensures such a $\delta$ exists). Then $|\Phi^{\prime}(p)|\leq L_{\delta}(\epsilon):=\epsilon/\sqrt{\delta(1-\delta)}$ for every subsequent $p$.
Step 3 – contraction. Using a gradient step $p_{n}=p_{n-1}+\eta\,\Phi^{\prime}(p_{n-1})$ with $\eta\in(0,1/L_{\delta}(\epsilon))$ gives
$|p_{n}-p_{n-1}|\leq\eta\,L_{\delta}(\epsilon)\,|p_{n-1}-p_{n-2}|=:\lambda\,|p_{n-1}-p_{n-2}|,\quad\lambda<1.$ (16)
Thus $h_{\epsilon}$ is a strict contraction on $([\delta,1-\delta],|\cdot|)$.
Step 4 – Banach fixed-point theorem. Because the metric space $\bigl{(}[\delta,1-\delta],\,|\cdot|\bigr{)}$ is complete and $h_{\epsilon}$ is a strict contraction on it, Banach’s fixed-point theorem guarantees
  1. (a)
    the existence and uniqueness of a fixed point $p^{\star}\in(\delta,1-\delta)\subset(0,1)$ such that $p^{\star}=h_{\epsilon}(p^{\star})$;
  2. (b)
    geometric convergence of the iterates generated by $p_{n}=h_{\epsilon}(p_{n-1})$:
    $|p_{n}-p^{\star}|\;\leq\;\lambda^{\,n}\,|p_{0}-p^{\star}|,\qquad 0<\lambda<1.$ (17)
Step 5 – bounded negative-advantage gradients. For every $\hat{A}_{i}<0$
$f_{\epsilon}^{\mathrm{RIFT}}(\rho_{i},\hat{A}_{i})=\max\!\bigl{(}\min(\rho_{i}\hat{A}_{i},\rho^{-}\hat{A}_{i}),\,c\hat{A}_{i}\bigr{)}\geq c\hat{A}_{i},$ (18)
so $|\partial f_{\epsilon}^{\mathrm{RIFT}}/\partial\rho_{i}|\leq|\hat{A}_{i}|$, preventing gradient explosions and guaranteeing stability. ∎

### v3

@@ 5.4 Ablation Study (Q3)
Building on the design choices introduced in Section˜4.2, we systematically ablate five components of RIFT: weighting scheme (Old-Weight vs. Equal-Weight), fine-tuning module (Scoring Head vs. All Head), KL regularization (w/ KL vs. w/o KL), policy clipping (Dual-Clip vs. PPO-Clip), and style preference (Normal vs. Aggressive). All experiments share identical settings, and results are reported in Table˜3.
Equal-Weight vs. Old-Weight. Replacing old-policy weighting with equal weighting eliminates the likelihood bias toward frequent modes and enables balanced updates across all candidates. This leads to improved exploitation of high-return rollouts and better multimodality preservation.
Scoring Head vs. All Head. Freezing the generation head is crucial for retaining trajectory-level realism and route-level controllability. Fine-tuning all heads (w/ All Head) disrupts the pre-trained generation head and slightly degrades realism metrics, whereas fine-tuning only the scoring head achieves better controllability without compromising realism.
w/ KL vs. w/o KL. Anchoring the scoring head to the IL pre-trained reference via KL regularization (w/ KL) constrains adaptation to a biased reference under asymmetric covariate shift. Removing this term improves controllability while maintaining realism, confirming that free adaptation of the scoring head yields more effective policy improvement.
Dual-Clip vs. PPO-Clip. Replacing dual-clip with standard PPO clipping (w/ PPO-Clip) results in overly conservative behaviors and reduced efficiency, as extreme negative updates can dominate and suppress positive learning signals. Dual-clip bounds such updates while preserving responsiveness to high-return rollouts, producing more realistic and efficient behavior.
Normal vs. Aggressive. Adopting a more aggressive reward that emphasizes efficiency increases route progress but also raises collision and off-road rates, illustrating the efficiency–safety trade-off. These results demonstrate that RIFT supports flexible style shaping while maintaining stability and multimodality. Additional qualitative insights on controllability are provided in Section˜D.1.
@@ 6 Conclusion
In this work, we propose a dual-stage AV-centric simulation framework that conducts IL pre-training in a data-driven simulator to capture trajectory-level realism and route-level controllability, followed by RL fine-tuning in a physics-based simulator to address covariate shift and enhance style-level controllability. During fine-tuning, we introduce RIFT, a novel group-relative RL fine-tuning strategy that evaluates all candidate modalities using the group-relative formulation combined with a surrogate objective for optimization, thereby enhancing style-level controllability and mitigating covariate shift, while preserving the trajectory-level realism and route-level controllability established in IL pre-training. Extensive experiments demonstrate that RIFT generates scenarios with superior realism and controllability, effectively revealing the limitations of modern AV systems and further bridging the gap between traffic simulation and reliable closed-loop evaluation. Due to space limitations, more discussion on limitations and future direction can be found in Appendix E.2.
@@ Appendix A Theoretical Analysis
A.1 Setting
For each $s\!\sim\!\mathcal{D}$, a frozen trajectory generation head yields $\mathcal{C}(s)=\{\tau_{i}\}_{i=1}^{G}$. The trajectory score head defines $\pi_{\theta}(\tau_{i}\mid s)$ on $\mathcal{C}(s)$. Finite-horizon simulation provides returns
$R_{i}(s)=\sum_{t=0}^{T}\gamma^{t}\,\mathrm{StateWiseRM}(\widetilde{\tau}_{i}^{t},s).$ (10)
Uniform (within-group) moments:
$\mu_{\mathrm{uni}}(s)=\tfrac{1}{G}\sum_{j=1}^{G}R_{j}(s),\qquad\sigma^{2}_{\mathrm{uni}}(s)=\tfrac{1}{G}\sum_{j=1}^{G}(R_{j}(s)-\mu_{\mathrm{uni}}(s))^{2}.$ (11)
Uniform, centered advantages:
$\hat{A}_{i}(s)=\frac{R_{i}(s)-\mu_{\mathrm{uni}}(s)}{\sqrt{\sigma_{\mathrm{uni}}^{2}(s)+\varepsilon}},\quad\frac{1}{G}\sum_{i=1}^{G}\hat{A}_{i}(s)=0.$ (12)
Let $\rho_{i}(\theta)=\pi_{\theta}(\tau_{i}\mid s)/\pi_{\theta_{\mathrm{old}}}(\tau_{i}\mid s)$. Define the RIFT surrogate
$\mathcal{J}_{\mathrm{RIFT}}(\theta)=\mathbb{E}_{s}\!\left[\frac{1}{G}\sum_{i=1}^{G}\psi\big(\rho_{i}(\theta),\hat{A}_{i}(s)\big)\right],$ (13)
with dual-clip kernel
$\psi(\rho,\hat{A})=\begin{cases}\min\!\big(\rho\,\hat{A},~\mathrm{clip}(\rho,1-\epsilon,1+\epsilon)\,\hat{A}\big),&\hat{A}\geq 0,\\[3.0pt] \max\!\Big(\min\!\big(\rho\,\hat{A},~\mathrm{clip}(\rho,1-\epsilon,1+\epsilon)\,\hat{A}\big),~c\,\hat{A}\Big),&\hat{A}<0,\end{cases}\quad(\epsilon>0,~c>1).$ (14)
Assumptions.
(A) Support floor: there exists $\pi_{\min}>0$ such that $\pi_{\theta_{\mathrm{old}}}(\tau_{i}\mid s)\geq\pi_{\min}$ for all $(s,i)$, and $\pi_{\theta}>0\Rightarrow\pi_{\theta_{\mathrm{old}}}>0$ on $\mathcal{C}(s)$. (B) Boundedness: $|\hat{A}_{i}(s)|\leq A_{\max}$. (C) Regularity: $\log\pi_{\theta}(\tau_{i}\mid s)$ is $L$-Lipschitz and $C^{2}$ on compact $\Theta$.
A.2 Listwise View and Diversity Pressure
Consider the unclipped uniform surrogate
$L_{\mathrm{RIFT}}(\theta)=\mathbb{E}_{s}\!\left[\frac{1}{G}\sum_{i=1}^{G}\rho_{i}(\theta)\,\hat{A}_{i}(s)\right]=\mathbb{E}_{s}\!\left[\frac{1}{G}\sum_{i=1}^{G}\frac{\hat{A}_{i}(s)}{\pi_{\theta_{\mathrm{old}}}(\tau_{i}\mid s)}\,\pi_{\theta}(\tau_{i}\mid s)\right].$ (15)
Proposition A.1 (Pairwise ascent and diversity).
Fix $s$ and shift an infinitesimal mass $\delta$ from $j$ to $i$ in $\pi_{\theta}(\cdot\mid s)$. Then $\delta L_{\mathrm{RIFT}}(\theta)=\tfrac{\delta}{G}\!\left(\tfrac{\hat{A}_{i}(s)}{\pi_{\theta_{\mathrm{old}}}(\tau_{i}\mid s)}-\tfrac{\hat{A}_{j}(s)}{\pi_{\theta_{\mathrm{old}}}(\tau_{j}\mid s)}\right)$. Hence ascent moves mass toward larger $\hat{A}/\pi_{\mathrm{old}}$, amplifying underrepresented high-quality candidates when $\pi_{\mathrm{old}}$ is peaky.
Corollary A.2 (Top-1 Fisher consistency under uniform reference).
If $\pi_{\theta_{\mathrm{old}}}$ is uniform on $\mathcal{C}(s)$ and $i^{\star}(s)=\arg\max_{i}\hat{A}_{i}(s)$ is unique, any global maximizer of $L_{\mathrm{RIFT}}$ concentrates $\pi_{\theta}(\cdot\mid s)$ on $i^{\star}(s)$.
A.3 Clipping as Stability Control
Clipping is a pointwise pessimistic transform: for any $x=\rho\,\hat{A}$,
$\min(\rho\,\hat{A},\mathrm{clip}(\rho)\hat{A})\leq\rho\,\hat{A}\$.
Summed over mixed signs, there is no global monotone lower bound for $L_{\mathrm{RIFT}}$; instead, clipping serves to bound the value and the gradient.
Lemma A.3 (Bounded values and gradients).
If $|\hat{A}|\leq A_{\max}$, then for all $(s,i)$: (i) Value bounds: $\psi\!\in[0,(1+\epsilon)\hat{A}$] for $\hat{A}\!\geq\!0$, and $\psi\!\in[c\,\hat{A},0]$ for $\hat{A}\!<\!0$. (ii) Gradient bounds:
$\bigg|\frac{\partial\psi}{\partial\log\pi_{\theta}}\bigg|\ \leq\ \begin{cases}(1+\epsilon)|\hat{A}|,&\hat{A}\geq 0,\\ c\,|\hat{A}|,&\hat{A}<0,\end{cases}$
and on the negative branch when the dual-clip is active ($\psi=c\,\hat{A}$) the derivative is 0.
A.4 Smoothness w.r.t. Policy Divergence
Write $w_{i}(s)=\hat{A}_{i}(s)/\pi_{\theta_{\mathrm{old}}}(\tau_{i}\mid s)$ and note $|w_{i}|\leq A_{\max}/\pi_{\min}$ under Assumption A with $\pi_{\min}=\inf_{s,i}\pi_{\theta_{\mathrm{old}}}(i\mid s)>0$ (label-smoothing in practice). Then the unclipped surrogate is linear in $\pi_{\theta}$:
$L_{\mathrm{RIFT}}(\theta)-L_{\mathrm{RIFT}}(\theta^{\prime})=\mathbb{E}_{s}\!\left[\frac{1}{G}\sum_{i}w_{i}(s)\big(\pi_{\theta}(i\mid s)-\pi_{\theta^{\prime}}(i\mid s)\big)\right].$
Lemma A.4 (Lipschitz continuity via KL).
For any $\theta,\theta^{\prime}$,
$\big|L_{\mathrm{RIFT}}(\theta)-L_{\mathrm{RIFT}}(\theta^{\prime})\big|\ \leq\ \frac{A_{\max}}{\pi_{\min}}\ \sqrt{2\,\mathbb{E}_{s}\!\big[\mathrm{KL}\big(\pi_{\theta}(\cdot\mid s)\,\|\,\pi_{\theta^{\prime}}(\cdot\mid s)\big)\big]}\,.$
Proof.
By Hölder and Pinsker: $|\sum_{i}w_{i}\Delta\pi|\leq\|w\|_{\infty}\,\|\Delta\pi\|_{1}\leq(A_{\max}/\pi_{\min})\sqrt{2\,\mathrm{KL}(\pi_{\theta}\|\pi_{\theta^{\prime}})}$, then average over $s$. ∎
Lemma A.5 (Lipschitz continuity of clipped surrogate).
Because $\partial\psi/\partial\pi_{\theta}(i|s)$ is bounded by $A_{\max}/\pi_{\min}$ whenever the active branch is differentiable,
$|\mathcal{J}_{\mathrm{RIFT}}(\theta)-\mathcal{J}_{\mathrm{RIFT}}(\theta^{\prime})|\ \leq\ \frac{A_{\max}}{\pi_{\min}}\ \sqrt{2\,\mathbb{E}_{s}\!\big[\mathrm{KL}\big(\pi_{\theta}(\cdot\mid s)\,\|\,\pi_{\theta^{\prime}}(\cdot\mid s)\big)\big]}.$
A.5 Variance and Enumeration
Let $f_{i}(s;\theta)=\psi(\rho_{i}(\theta),\hat{A}_{i}(s))$. Exact enumeration yields $\operatorname{Var}\!\big(\tfrac{1}{G}\sum_{i}f_{i}\,\big|\,s\big)=0$ (assuming $\widetilde{\tau}_{i}$ and their evaluations are fixed during the update; otherwise, environment randomness still induces nonzero variance), while sampling $i$ i.i.d. within the group gives conditional variance $\operatorname{Var}(f_{i}\mid s)/N$ for $N$ samples.
A.6 Convergence of Stochastic Ascent
Theorem A.6 (Convergence to a stationary point).
Under Assumptions A–C, with step sizes $\eta_{k}>0$, $\sum_{k}\eta_{k}=\infty$, $\sum_{k}\eta_{k}^{2}<\infty$, and unbiased bounded-variance stochastic subgradients, the iterates of stochastic subgradient ascent on $\mathcal{J}_{\mathrm{RIFT}}$ satisfy
$\liminf_{k\to\infty}\mathbb{E}\!\big[\operatorname{dist}\big(0,\partial^{C}\mathcal{J}_{\mathrm{RIFT}}(\theta_{k})\big)\big]=0,$
where $\partial^{C}$ denotes the Clarke generalized gradient.
Sketch.
By Lemma A.3, generalized gradients are uniformly bounded; regularity of $\log\pi_{\theta}$ on compact $\Theta$ implies Lipschitz continuity. Robbins–Monro / Kushner–Yin results for non-smooth stochastic approximation apply. ∎
A.7 Why RIFT Preserves Multimodality
By Proposition A.1, ascent compares $\hat{A}/\pi_{\mathrm{old}}$: under peaky $\pi_{\mathrm{old}}$, underrepresented high-$\hat{A}$ candidates receive stronger positive updates, preserving and enhancing diversity. In the special case $\pi_{\mathrm{old}}$ is (approximately) uniform, RIFTreduces to a listwise ranking ascent that directly promotes larger $\hat{A}$.